OpenStudy (anonymous):

Find the perimeter of a quadrilateral with vertices at L (−4, 0), M (4, 6), N (7, 2), O (−1, −4). Round your answer to the nearest hundredth when necessary. 30 units 42 units 48 units 50 units

OpenStudy (zpupster):

plot points N M L O then use distance formula between points d=sqrt((x2-x1)^2+(y2-y1)^2) i will do it between L and M M(4,6) or x2,y2 L(-4,0) or x1,y1 plug them in the formula d = sqrt(4-(-4)^2 + (6-0)^2 d= sqrt 8^2 + 6^2 d=sqrt(64+36) d=sqrt100 d=10 this is one side, you will have 3 more sides to obtain then add all the lengths of the 4 sides together and get the perimeter